OpenStudy (anonymous):

tim spent $24 at the grocery store. this was six dollars less than than two times what he spent at the game store . how much did he spend at the game store? A. $15 B. $25 C. $35 D. $45

OpenStudy (anonymous):

2x-6=24 x=24+6/2 =30/2 =15$ hope its correct

OpenStudy (anonymous):

unfortunatley 9 is not an answer:/

OpenStudy (anonymous):

its 15$ :D :P that was a miscalculation simply :D

OpenStudy (anonymous):

$15 24 + 6= 30 then you divide it by two to get the answer 30/2 = 15

OpenStudy (anonymous):

THANK YOU(:<33

OpenStudy (anonymous):

your welcome can i have a medal

OpenStudy (anonymous):

:D Thanks :D You're welcome and @loser123456789876 Happy :D ??

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

you happy